OpenStudy (anonymous):

Need some help for this.. A 10-foot ladder leans against a wall with its foot braced 3 feet from wall’s base. How far up the wall does the ladder reach? Show your work. Round answer to the nearest hundredths.

Mehek (mehek14):

 do you know the pythagorean theorem?

OpenStudy (anonymous):

a^2+ b^2 = c

Mehek (mehek14):

 \(a^2+b^2=c^2\)

Mehek (mehek14):

you have a and c \(3^2+b^2=10^2\)

Mehek (mehek14):

tell me what \(3^2=\\10^2=\)

OpenStudy (anonymous):

9 and 100

Mehek (mehek14):

ok subtract 100-9

OpenStudy (anonymous):

91

Mehek (mehek14):

\(\sqrt{91}\)

OpenStudy (anonymous):

Oh. can't make a radical symbol.

OpenStudy (anonymous):

\[10 \sin \left(\cos ^{-1}\left(\frac{3}{10}\right)\right)=\sqrt{91} \]

OpenStudy (anonymous):

Okay thanks @Mehek14

Mehek (mehek14):

and then just round your answer

OpenStudy (anonymous):

How do i round √91?

Mehek (mehek14):

well if you put √91 in the calculator, you get 9.53939201417 round that to 9.54

OpenStudy (anonymous):

Oh didn't put it in a calculator thanks for telling me that XwX

OpenStudy (anonymous):

So i am finish @Mehek14?

Mehek (mehek14):

yes 9.54 is the final answer

OpenStudy (anonymous):

Okay thanks @Mehek14

Mehek (mehek14):

yw